a massless string passes over a frictionless pulley and carries two masses m1 and m2(m1>m2) if g is acc. due to gravity then thrust on the pulley is?


 

Grade:10

1 Answers

vikas askiitian expert
13 years ago

at mass m1

          m1g - T = m1a ...........1

at mass m2

                 T-m2g = m2a ..............2

solving these 

         T= 2m1m2/(m1+m2) ..........3

net force on the pulley = sum of tension in both the strings

                 F= T+T=2T

                F(thrust)  =4m1m2/(m1+m2) .........4